OpenStudy (anonymous):

Find the area of a triangle whose base is 5cm and whose height is 12cm.

OpenStudy (anonymous):

Area = (b*h)/2

OpenStudy (anonymous):

\[\frac{ 5*12 }{ 2 }=30\]

OpenStudy (anonymous):

thanks I was forgetting to divide :/